OpenStudy (anonymous):

perpendicular to the line y=3x-3; containing the points (2,-1)

OpenStudy (mathmate):

@j24rodriguez can you tell me the slope of the given line?

OpenStudy (uri):

So, the slope of the given line is 3 since it is in y=mx+b form. Thus, the perpendicular line has slope m = -1/3.

OpenStudy (anonymous):

i believe it's 4 if i'm not mistaken

OpenStudy (mathmate):

@j24rodriquez Compare y=3x-3 with y=mx+c m is the slope, c is a constant. Can you find the slope now?